mikewwe13:

The number of students enrolled in a new course as a function of time can be represented by the function. f(x) = (4)^x − 1 What is the average increase in the number of students enrolled per hour between hours 2 and 4? Enter your answer in the box. _______________

mikewwe13:

@Vocaloid

Mehek:

First you'll plug 2 and 4 for x to find f(x) and then you'll find the rate of change

Mehek:

f(2) = 4^2 - 1 Can you simplify 4^2 - 1?

mikewwe13:

it is 15

Mehek:

Correct so when x = 2, f(x) = 15 Now for x = 4 f(4) = 4^4 - 1 Could you simplify that

mikewwe13:

255

Mehek:

Correct Now the average is \(\bf\dfrac{f(4)-f(2)}{4-2}=\dfrac{255-15}{4-2}=?\)

mikewwe13:

120

Mehek:

So 120 is your answer

mikewwe13:

ok thx for that and could you help me with other questions if you have the time if it's ok with u?

Mehek:

Sure

mikewwe13:

ok hold on

mikewwe13:

Drag and drop the answer into the box to identify each function as linear, exponential, or neither. linear exponential neither

mikewwe13:

i'll send the pic

mikewwe13:

1 attachment
Mehek:

You'd have to find the rate of change Know how to?

mikewwe13:

well i think i have it

mikewwe13:

the first is exponential the second is linear correct me

Mehek:

I think you have it backwards lol For the first one it's a constant rate of 1.5 which is \(\bf linear\) And then for the second one, it seems like it's 2^x, so 2^0 = 1, 2^3 = 8, 2^6 = 64 which makes it \(\bf exponential\)

mikewwe13:

oh so i had but it's backwards

Mehek:

Yep

mikewwe13:

i'll send the pic

mikewwe13:

1 attachment
mikewwe13:

Function f is an exponential function. By what factor does the output value increase as each input value increases by 1? Enter your answer in the box. ____________

Mehek:

Any idea?

mikewwe13:

no

mikewwe13:

do you ?

Mehek:

I know the answer but am not sure how to explain it

Mehek:

I basically kept multiplying by 2 each time and got the same thing

1 attachment
Mehek:

Idk why there's an image

mikewwe13:

ok

mikewwe13:

so what is it ?

mikewwe13:

mehek ?